### v1

# Mapping the Global Election Landscape on Social Media in 2024

Abstract
In 2024, half of the global population is expected to participate in elections, offering researchers a unique opportunity to study online information diffusion and user behavior. This study investigates the media landscape on social media by analyzing Facebook posts from national political parties and major news agencies across Europe, Mexico, and India. Our methodology identifies key topics and evaluates public interaction, reflecting broader trends in political engagement. Using Principal Component Analysis, we distil these topics to uncover patterns of correlation and differentiation. This approach reveals dominant themes that engage global audiences, providing critical insights into the interplay between public opinion and digital narratives during a major electoral cycle. Our findings highlight how different topics resonate across political spectrums, shaping political debate and offering a comprehensive view of the interaction between media content, political ideology, and audience engagement.
Introduction
The advent and proliferation of social media platforms have fundamentally altered the information landscape and how we interact online [1], transforming these platforms into essential tools for information [2, 3], entertainment [4, 5], and personal communication [6]. As these platforms have become integrated into our daily lives, they have also merged entertainment-driven business models with complex social dynamics [7], raising significant concerns about their impact on social dynamics, particularly in terms of polarization [8, 9, 10, 11], where content tailored to reinforce existing beliefs may eventually end up fostering societal division [12] and misinformation spreading [13, 14, 15].
Thus, social media platforms have revealed to be critical arenas which saw an explosion of information and misinformation for global events, including the COVID-19 pandemic [16, 17], political events [18, 19, 20] and discussions on emerging technologies such as large language models [21, 22] and their implications for communication and automation. Moreover, these platforms play a pivotal role in consolidating biased views of the political landscape, potentially influencing public opinion and voter behavior during elections [23, 19] through the rapid dissemination and amplification of political content [24].
Research on how platform-specific effects influence social dynamics has highlighted several challenges [12, 25, 10, 26, 27]. Indeed, these interactions often exhibit persistent patterns despite different platforms, topics, and contexts, suggesting underlying consistencies in online human behavior [1]. Online users show the tendency to selectively expose to information [28, 29], preferring content that aligns with their pre-existing beliefs while avoiding contrary evidence [30, 31]. This behavior may fosters the emergence of homophilic communities, also know as echo-chambers [8], which significantly influence belief formation and communication methods [12], especially during delicate periods such as elections.
As the 2024 elections approach, with a substantial global population expected to vote, Facebook is a crucial platform for electoral campaigning. This period offers an invaluable opportunity to comprehensively analyze how different countries use social media for news consumption, public debate, and electoral influence.
In this study, we analyze content from news agencies and political parties on Facebook to see how it affects user interactions (including reactions, comments and shares) and examine variations between countries and the political leanings of parties. Using Principal Component Analysis, our analysis shows the main trends of the public discourse within the social media ecosystem, allowing us to identify patterns of both correlation and differences among key themes based on geographic and political distinctions.
Our comparative study across multiple countries aims to provide a multilateral understanding of how discussions on social media reflect electoral dynamics in a pivotal election year. Additionally, this paper explores how different countries, and thus socio-cultural factors, shape user interactions and interests on social media, contributing to the global dialogue on democracy and public discourse.
The structure of the paper is the following: We first assess the level of engagement on social media across various countries, then we explore the controversy surrounding discussions on diverse topics and, finally, we categorize the information and topics based on political leanings, providing a comprehensive overview of the digital electoral landscape in 2024.
Materials and methods
Dataset
For our analysis, we manually compiled detailed lists of Facebook pages from each country in the European Union, as well as the UK, India, and Mexico. These lists include pages linked to the websites of political parties represented in national parliaments and major news agencies in each country. We also included major American news agencies and public Facebook pages of influential politicians, identified using the YouGov public survey platform [32]. Our dataset includes 508 news agencies and 336 political parties across 31 countries, capturing a total of approximately 4.2 million and 176 thousand posts between September 1, 2023, and May 1, 2024. For each post, we recorded the page name, posting time, textual content, and aggregated metrics like the number of reactions and comments.
News outlet and parties labeling
In our study, we analyze Facebook pages from various countries that provide diverse perspectives on current events, ranging from the progressive left to the conservative right. We manually check political parties’ ideological leanings, since they are often clear and aligned with international coalitions, such as those within the European Parliament. Unlike political parties, news outlets usually have less explicit political alignments. To classify these, we use ratings from Media Bias/Fact Check (MBFC), an independent website that evaluates news agencies based on factuality and political bias. While MBFC recognizes a wide range of political leanings, for analytical simplicity, we consolidate these into three primary categories: ‘Left,’ ‘Center,’ and ‘Right.’
Process of Topic Modelling
We implemented a two-step process to assess the topics discussed in each post. We processed each country separately since each one had posts in their primary language. For each country, we selected a random sample of 50,000 posts made by news organizations. We then cleaned the text of these posts by removing links, hashtags, and non-alphanumeric characters, such as emojis. Subsequently, we applied the BERTopic clustering algorithm to perform topic clustering for each country. The resulting clusters were labelled with overarching themes such as ‘economy,’ ‘crime,’ ‘entertainment,’ and ongoing conflicts in Ukraine and the Middle East (identified as ‘mena’ in the analysis). Notice that, Finland and Sweden did not produce any discernible topics during the clustering phase. To infer the topic of all posts, we computed the term frequency-inverse document frequency (tf-idf) values of the words from the posts within each cluster, treating the concatenation of all posts from each macro-topic as a single document. This allowed us to identify the macro-topic that most closely aligned with the textual content of each post based on the tf-idf scores of the words. Fig. 1 presents the share of posts discussing each topic. As shown, political parties predominantly post about internal politics, whereas ‘lighter’ topics such as entertainment or sports garner significantly more attention from news outlets.
Fig 1: Topics discussed as share of total classified posts. The plurality of posts from political parties is about internal politics.
Principal component analysis
Principal Component Analysis (PCA) [33] is a method for linear dimensionality reduction, widely used to summarize and visualize the information of datasets containing a large number of individuals or observations in lower dimensions containing the greater possible variance [34, 35].
In this work, we mostly use biplots [36] and methods to visualize the overall relation of individuals or variables. Therefore we briefly recall some aspects used in this work.
Consider a matrix $M$ of dimension $n\times m$ in which rows are individuals and columns are variables measured on each individual. The entry $M_{i,j}$ contains the $j-$th variable of individual $i$.
After applying PCA to the columns of $M$ (each column is a variable) it is common to represent the results in a biplot, a two-dimensional space containing the larger possible variance for both variables and individuals, even if their coordinates are not constructed onto the same space.
Here, we summarize their properties:
  • the distance between the variables and the origin is a measure of how well that variable is explained in two dimensions;
  • if $i$ and $j$ are variables, we have $\cos\theta_{ij}=r_{ij}$, i.e. the cos of the angle between two variables is equal to their correlation. Therefore, positively correlated variables are grouped together, while negatively correlated variables are positioned on opposite quadrants;
  • individuals with similar values of variables are grouped together;
  • an individual that is on the same (opposite) side of a certain variable has a high (low) value for that variable.
Note that variables and individuals could also be represented separately, keeping the properties listed above.
In this work, we use the packages FactoMineR [37] and factoextra [38] to apply PCA and construct the relative biplots. In particular, we represent each individual as a black point, while variables are represented using red vectors starting from the origin.
Results and Discussion
Engagement across countries
We start the analysis by exploring the distribution of user engagement, which we measure through the reactions received by posts. Fig. 2 illustrates the distribution of these interactions for both news outlets and political parties.
Fig 2: Distribution of reactions. $(a)$ News outlets dataset and $(b)$ Parties dataset.
We observe that reactions across different types of content exhibit similar distributions. This pattern holds true for both news agencies and political parties, as illustrated in Fig. 2 (a) and (b). However, a significant deviation occurs in the likes received by political parties, where the mode of the distribution is approximately $10$. This indicates a reduced frequency of posts that receive very few likes for these entities. Moreover, consistent with prior research [1], heavy tails characterize all the observed distributions, suggesting that while most posts receive relatively few reactions, a small number accumulates a disproportionately large number.
Interestingly, the number of reactions received by posts from political parties and news agencies falls within similar orders of magnitude, with tails ranging between $10^{3}$ and $10^{5}$ reactions. Engagement distribution patterns for each country are displayed in Fig. S3 for news agencies and S4 for political parties, exhibiting the same consistent patterns.
Controversial topics across countries
In this section, we investigate which topics capture more contentious debate by utilizing a metric derived from the number of ‘love’ and ‘angry’ reactions each post receives. Specifically, we define the love-angry score [4] for post $i$, denoted as $LA(i)$, using the following equation:
$LA(i)=\frac{a_{i}-l_{i}}{a_{i}+l_{i}},$
where $a_{i}$ represents the number of angry reactions and $l_{i}$ the number of love reactions received by post $i$. This metric serves to quantify the degree of divisiveness elicited by each post. In particular, the score takes values in $[-1,+1]$, with $-1$ ($+1$) corresponding to posts that have received only love (angry) reactions. Fig. 3 displays the distribution of love-angry scores of news agencies and political parties grouped according to their political leaning.
Fig 3: Boxplots of love-angry scores received for each topic
The results show that news agencies often receive higher ‘love-angry’ scores on their posts than political parties. This suggests news agencies reach a wider audience, attracting diverse reactions. On the other hand, posts from political parties generally see lower scores, indicating they might be engaging a more specific group of followers that are likely to share similar views and which could imply less debate and more agreement, typical of an echo chamber effect [12, 8].
Topics in each country
In this section, we aim to highlight the most discussed topics in each country, using the total number of interactions as a proxy for their relative perceived importance.
We define $C$ as the number of distinct countries and $T$ as the number of topics discussed. To mitigate potential biases due to the varying numbers of news outlets across countries, we construct a matrix $M$ of dimensions $C\times T$, where the entry $M_{ij}$ represents the fraction of total engagement received in topic $j$ by country $i$.
To distinguish between the discussions led by news outlets and political parties, we create two separate matrices, $M^{news}$ and $M^{parties}$. Principal Component Analysis (PCA) is then applied to these matrices to derive a two-dimensional representation that explains the most variance; further details can be found in the Materials and Methods section. Fig. 4 presents the biplots from $M^{news}$ and $M^{parties}$.
Fig 4: Most engaged topics in each country, as revealed by PCA.
Notably, news outlets in India demonstrate unusually high engagement compared with other countries on topics such as the Middle East (abbreviated as MENA) and Ukraine. For most other countries, engagement patterns on the first principal component, which captures the largest variance, are relatively uniform, suggesting similar interest levels in entertainment, crime, and politics.
However, specific topics such as education, religion, technology, abortion, and africa show an inverse relationship with others like climate, health, protest, and economy. Countries like Ireland, Austria, Malta, and Spain are more engaged with the latter topics while less so with the former. From the perspective of political parties, engagement patterns vary significantly. For instance, parties in Northern European countries like Denmark, Belgium, and Latvia actively engage more with climate change, economy, education, health, abortion, technology, and sports. In contrast, topics such as protest, asia, religion, and ukraine obtain higher interactions in countries like Italy, India, Hungary, and Slovakia. This highlights a clear divergence in the focus of discussions between news outlets and political entities.
Topics and political leaning
In the previous section, we extracted the most engaging topics from each country without considering the specific landscape of news outlets and political parties within those countries. In this section, we address this gap by applying a similar analysis distinctly for each nation, incorporating an additional dimension of political leaning (details are provided in the Materials and Methods section).
We focus on the news outlets case (the parties are treated similarly). Following a method similar to the previous one, we consider a country $c$ with $n_{c}$ labeled news outlets discussing $m_{c}$ topics. For each country, we construct a matrix $M^{c}$ of dimension $n_{c}\times m_{c}$, where $M^{c}_{ij}$ represents the fraction of engagement received on topic $j$ by news outlet $i$. Fig. 5 shows the biplot resulting after applying PCA to $M^{c}$ on each country having $n_{c}\geq 3$.
Fig 5: Biplot of news outlets grouped by engaged topics for countries with $n_{c}\geq 3$.
We conduct a similar analysis for political parties, whose results are displayed in Fig. S1 for all countries having $n_{c}\geq 3$.
Interestingly, the analysis reveals that for both news outlets and political parties, the political leaning of each entity does not significantly influence how engagement is distributed across topics. This suggests that no topics are exclusively engaged by entities with a specific political leaning.
One possible explanation for the observed engagement patterns across different political leanings can be attributed to the overall narrative style of each page. Even when discussing the same topics, different pages may present perspectives that resonate more closely with specific audiences. However, due to the limitations of our classification approach, PCA primarily captures the general content themes without distinguishing between nuanced perspectives.
Despite this limitation and the potential imbalances in labeling across different countries, our findings suggest that news outlets and political parties, regardless of their political orientation —left, center, or right— are likely to engage with similar topics.
To delve deeper into the engagement dynamics specific to each political leaning, we employ a similar analytical procedure. In particular, we construct a matrix $M^{b}$ in which the rows are countries and columns are topics. Differently from before, the generic entry $M^{b}_{i,j}$ is the fraction of engagement received by news outlets of country $i$, having political leaning $b\in\{left,right\}$, in topic $j$. PCA is then applied to these matrices to yield a two-dimensional representation of the topic discussed by a specific political leaning in each country. The results are reported in Fig. 6.
Fig 6: Most engaging topics of news outlets and parties if only a specific leaning is considered.
We observe that, for what concerns left news outlets Poland, Spain, USA, Ireland and UK are getting high engagements in topic such as abort, religion, climate and tech. At the same time Netherlands, France and Germany are more focused on education, economy, africa and migration. On the other hand, right news outlet in Italy, Germany, UK and Netherlands are getting high engagements in topics such as gender, migration and ukraine. Interestingly, USA, Portugal, Denmark and Spain are more interested in religion, abortion, sport and crime. Finally, only right news outlets from France and Austria seems to get engagement with the climate topic.
We observe a more heterogeneous scenario for political parties, where all topics are treated more homogeneously.
Conclusion
This study analyzes how different political leanings and media outlets engage audiences on social media, revealing patterns in how various topics resonate with people globally. Despite challenges with data classification and PCA’s limitations, our findings show clear trends in topic engagement that go beyond political lines. Our analysis indicates that while news outlets and political parties often discuss similar broad topics, the intensity and nature of this engagement differ based on political orientation. Specifically, right-leaning outlets and parties focus more on politics, religion, and migration. In contrast, left-leaning groups emphasize education, health, and technology. This suggests that the left may be talking about things that need to capture more people’s interest than the topics the right focuses on. The PCA biplots show differences in topic engagement, but there isn’t a sharp divide between the left and right. This means that even politically polarized groups can find common ground on specific issues from different perspectives. For instance, both sides might discuss the economy but approach it from their unique viewpoints. These findings are important for policymakers, media strategists, and researchers. Understanding these engagement patterns can create more effective messages that appeal to a wider audience, bridging divides and fostering better public discourse. This study also offers practical insights for designing social media campaigns considering ideological differences and audience preferences.
Acknowledgments
Supporting information
Fig S1: Biplot of parties grouped by engaged topics.
Fig S2: More engaging topics for parties.
Fig S3: Reaction distributions to news posts across countries.
Fig S4: Reaction distributions to posts from political parties across countries.